# Targeted Data Fusion for Causal Survival Analysis Under Distribution Shift

### v1

@@ D.2.3 Lemmata for the source-site estimator
by Bayes’s rule. Furthermore, by Duhamel equation in Gill & Johansen (1990) used in Lemma D.5 again, $\mathbb{P}[\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P}_{\infty})]-\theta^{0}(t,a)$ equals
$\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}S^{k}_{\infty}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})}{S^{k}_{\infty}(y\mid a,\mathbf{X})}\left\{\frac{\omega^{k,0}_{\infty}(\mathbf{X})G^{k}(y\mid a,\mathbf{X})\pi^{k}(a\mid\mathbf{X})}{\omega^{k,0}(\mathbf{X})G^{k}_{\infty}(y\mid a,\mathbf{X})\pi^{k}_{\infty}(a\mid\mathbf{X})}-1\right\}(\Lambda^{k}_{\infty}-\Lambda^{k})(dy\mid a,\mathbf{X})\right].$ (7)
@@
By (6) with $\pi_{\infty}^{k}=\pi^{k}$, $\omega^{k,0}_{\infty}=\omega^{k,0}$, $G_{\infty}^{k}=G^{k}$, and $S_{\infty}^{k}=S^{k}$,
$\widehat{\theta}^{k,0}_{n}(t,a)-\theta^{0}(t,a)$ $=\mathbb{P}_{n}[\varphi^{*k,0}_{t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{k,0}_{n,m,t,a}-\varphi^{k,0}_{t,a}\right]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{k,0}_{t,a}-\theta^{0}(t,a)\right].$
By Conditions D.1 and D.2, the second summand on the right-hand-side is $o_{p}(n^{-1/2})$ by Lemma D.10. By Lemma D.11, $\mathbb{P}[\widehat{\varphi}^{k,0}_{t,a}]-\theta^{0}(t,a)$ equals
$\mathbb{E}\left[\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\widehat{S}^{k}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})}{\widehat{S}^{k}_{m}(y\mid a,\mathbf{X})}\left\{\frac{\widehat{\omega}^{k,0}_{m}(\mathbf{X})G^{k}(y\mid a,\mathbf{X})\pi^{k}(a\mid\mathbf{X})}{\omega^{k,0}(\mathbf{X})\widehat{G}^{k}_{m}(y\mid a,\mathbf{X})\widehat{\pi}^{k}_{m}(a\mid\mathbf{X})}-1\right\}(\widehat{\Lambda}^{k}_{m}-\Lambda^{k})(dy\mid a,\mathbf{X})\right].$
By a similar decomposition in the proof of Theorem 2.9 in Section D.1.3 using Duhamel equation in Gill & Johansen (1990), and by notation in Condition D.4, we find that the above bias term can be bounded by $\eta^{2}\{r_{n,t,a,1}^{k}+r_{n,t,a,2}^{k}+r_{n,t,a,3}^{k}\}$ over $m$. Since $M^{-1}\sum_{m=1}^{M}n^{-1}Mn_{m}\leq 2$, we have
@@
by Condition D.4. This established the pointwise RAL property: $\widehat{\theta}^{k,0}_{n}(t,a)=\theta^{0}(t,a)+\mathbb{P}_{n}(\varphi^{*k,0}_{t,a})+o_{p}(n^{-1/2})$. Since $\varphi^{*k,0}_{t,a}$ is uniformly bounded, $\mathbb{P}\{(\varphi^{*k,0}_{t,a})^{2}\}<\infty$ and since $\mathbb{P}\{\varphi^{*k,0}_{t,a}\}=0$, it follows that
@@
For the uniform RAL, by conditions D.1, D.2 and D.4 and Lemma D.10,
@@
which is $o_{p}(n^{-1/2})$ by Condition D.5. Therefore, $\sup_{u\in[0,t]}\left|\widehat{\theta}^{k}_{n}(u,a)-\theta^{0}(u,a)-\mathbb{P}_{n}(\varphi^{*k,0}_{u,a})\right|=o_{p}(n^{-1/2})$. Since $\{\varphi^{*k,0}_{u,a}:u\in[0,t]\}$ is a uniformly bounded $\mathbb{P}$-Donsker class by Lemma D.9, $\{n^{1/2}\mathbb{P}_{n}\{\varphi^{*k,0}_{u,a}\}:u\in[0,t]\}$ converges weakly to a tight mean-zero Gaussian process with covariance $(u,v)\mapsto\mathbb{P}(\varphi^{*k,0}_{u,a}\varphi^{*k,0}_{v,a})$. ∎
@@
If we only need the consistency of $\widehat{\theta}_{n}^{k}(t,a)$, then condition $\pi_{\infty}^{k}=\pi^{k}$, $\omega^{k,0}_{\infty}=\omega^{k,0}$, $G_{\infty}^{k}=G^{k}$, and $S_{\infty}^{k}=S^{k}$ can be replaced by the following statement: For $\mathbb{P}$-almost all $\mathbf{x}$, there exist measurable sets $\mathcal{S}^{k}_{x},\mathcal{G}^{k}_{x}\subseteq[0,t]$ such that $\mathcal{S}^{k}_{x}\cup\mathcal{G}^{k}_{x}=[0,t]$ and $\Lambda^{k}(u\mid a,\mathbf{x})=\Lambda_{\infty}^{k}(u\mid a,\mathbf{x})$ for all $u\in S^{k}_{x}$ and $G(u\mid a,\mathbf{x})=G^{k}_{\infty}(u\mid a,\mathbf{x})$ for all $u\in\mathcal{G}^{k}_{x}$. In addition, if $\mathcal{S}^{k}_{x}$ is a strict subset of $[0,t]$, then $\pi^{k}(a\mid\mathbf{x})=\pi^{k}_{\infty}(a\mid\mathbf{x})$ and $\omega^{k,0}(\mathbf{x})=\omega^{k,0}_{\infty}(\mathbf{x})$ as well. Then, $\widehat{\theta}_{n}^{k}(t,a)$ is consistent if Conditions D.1 and D.2 hold, and it is uniform consistent if Condition D.3 also holds. This statement could be interpreted as that at a given time $t$, if either (i) the conditional survival model $S^{k}$; or (ii) all other nuisance functions $G^{k}$, $\pi^{k}$ and $\omega^{k,0}$ are correctly specified (with other conditions above), $\widehat{\theta}_{n}^{k}(t,a)$ is consistent.
The proof sketch of Remark D.12 is similar to that under Remark D.6 via examining the bias term (7), which is omitted here.
@@ D.3 Theory for the federated estimator
Let the set of all source site indices be $\mathcal{S}={1,\dots,K-1}$. We define the oracle selection space for $\boldsymbol{\eta}_{t,a}$ as:
$\mathcal{S}^{*}_{t,a}=\{k\in\mathcal{S}:\theta^{k}(t,a)=\theta^{0}(t,a)\},~{}~{}\mathbb{R}^{S^{*}_{t,a}}=\{\boldsymbol{\eta}_{t,a}\in\mathbb{R}^{K-1}:\eta^{j}_{t,a}=0,\forall j\not\in\mathcal{S}^{*}_{t,a}\}.$
This space is both time- and treatment-varying, indicating that a source site may not consistently be useful or unhelpful across different time points or treatments. However, it offers the advantage of increased flexibility and adaptivity, allowing for more effective borrowing of information at different points along the survival functions. Based on the theory presented in Section D.2, for $k\in\mathcal{S}^{*}_{t,a}$, the site-specific estimator $\widehat{\theta}^{k,0}_{n}(t,a)$ is consistent for $\theta^{0}(t,a)$ for any given $t\in[0,\tau]$ and $a\in\{0,1\}$.
We begin by assuming fixed $\boldsymbol{\eta}_{t,a}=(\eta_{t,a}^{0},\eta_{t,a}^{1},\dots,\eta_{t,a}^{K-1})$. We invoke Lemmata 4 and 5 in Han et al. (2021), which state that the proposed adaptive estimation for $\eta_{t,a}^{k}$ as shown in (2.3.2) allows for (i) the recovery of the optimal $\bar{\eta}^{k}_{t,a}$ by the estimator $\widehat{\eta}^{k}_{t,a}$, and (ii) the uncertainty induced by $\widehat{\eta}^{k}_{t,a}$ is negligible when estimating $\theta^{0}(t,a)$. We require regularity Conditions D.1, D.2 and D.4 for the pointwise convergence result in Theorem D.7 hold. Let us denote the federated estimator by plugging-in the fixed $\boldsymbol{\eta}_{t,a}$ as
@@
Let us write
$\xi^{0,(1)}(\mathcal{O})$ $=S^{0}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\pi^{0}(a\mid\mathbf{X})}\bigg{\{}\frac{\mathbb{I}(Y\leq t,\delta=1)}{S^{0}(Y\mid a,\mathbf{X})G^{0}(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\Lambda^{0}(du\mid a,\mathbf{X})}{S^{0}(u\mid a,\mathbf{X})G^{0}(u\mid a,\mathbf{X})}\bigg{\}},$
$\xi^{k,0,(1)}(\mathcal{O})$ $=\omega^{k,0}(\mathbf{X})S^{k}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}\bigg{\{}\frac{\mathbb{I}(Y\leq t,\delta=1)}{S^{k}(Y\mid a,\mathbf{X})G^{k}(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\Lambda^{k}(du\mid a,\mathbf{X})}{S^{k}(u\mid a,\mathbf{X})G^{k}(u\mid a,\mathbf{X})}\bigg{\}},$
@@
$\widehat{\theta}^{\text{fed}}_{n}(t,a;\boldsymbol{\eta}_{t,a})-\theta^{0}(t,a)$ $=\left(1-\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\right)\left\{\widehat{\theta}^{0}_{n}(t,a)-\theta^{0}(t,a)\right\}+\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\left\{\widehat{\theta}^{k,0}_{n}(t,a)-\theta^{0}(t,a)\right\}$
@@
The asymptotic variance of $\widehat{\theta}^{\text{fed}}_{n}(t,a;\boldsymbol{\eta}_{t,a})$ equals the variance of the influence function of (D.3). Let us denote it as $\mathcal{V}(\boldsymbol{\eta}_{t,a})$. Under the assumption of i.i.d. participants within each site, we have
$\mathcal{V}(\boldsymbol{\eta}_{t,a})$ $=\left(1-\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\right)^{2}\frac{\mathbb{V}\{\xi^{0,(2)}(\mathcal{O}_{i})-\xi^{0,(1)}(\mathcal{O}_{i})\mid R_{i}=0\}}{\mathbb{P}(R_{i}=0)}$
@@
$\sqrt{n}\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a;\boldsymbol{\eta}_{t,a})-\theta^{0}(t,a)\right\}\to_{d}\mathcal{N}(0,\mathcal{V}(\boldsymbol{\eta}_{t,a})).$
@@
$\bar{\boldsymbol{\eta}}_{t,a}=\arg\min_{\eta^{k}_{t,a}=0,\forall k\not\in\mathcal{S}^{*}_{t,a}}\mathcal{V}(\boldsymbol{\eta}_{t,a}).$ (10)
By leveraging Lemmata 4 and 5 from Han et al. (2021), we can recover the optimal weights $\bar{\boldsymbol{\eta}}_{t,a}$ with negligible uncertainty for estimating $\theta^{0}(t,a)$ if we estimate ${\boldsymbol{\eta}}_{t,a}$ using (2.3.2), akin to adaptive Lasso (Zou, 2006; Fan et al., 2024). The consistency of $\widehat{\mathcal{V}}(\widehat{\boldsymbol{\eta}}_{t,a})$ follows when we can effectively approximate ${\mathcal{V}}(\bar{\boldsymbol{\eta}}_{t,a})$ with $\widehat{\mathcal{V}}(\widehat{\boldsymbol{\eta}}_{t,a})$. Thus,
$\sqrt{n/\widehat{\mathcal{V}}(\widehat{\boldsymbol{\eta}}_{t,a})}\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a)-\theta^{0}(t,a)\right\}\to_{d}\mathcal{N}(0,1).$
@@
To establish that the asymptotic variance of $\widehat{\theta}^{\text{fed}}_{n}(t,a;\widehat{\boldsymbol{\eta}}_{t,a})$ is strictly smaller than that of the estimator based solely on the target data $\widehat{\theta}^{0}_{n}(t,a)$, we adopt Proposition 1 in Han et al. (2021) with a modified informative source condition (modified Assumption 3(b) in Han et al. (2021)).
Specifically, for each source site $s\in\mathcal{S}^{*}_{t,a}$, we define $\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})$ a federated estimator where $\eta^{s}_{t,a}$ is the optimal ensemble weight of site $s$ if we only consider target site and this source site $s$ for the federation. Then, the modified informative source condition is $\left|\text{Cov}\left[\sqrt{n}\widehat{\theta}^{0}_{n}(t,a),\sqrt{n}\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})-\widehat{\theta}^{0}_{n}(t,a)\right\}\right]\right|\geq\varepsilon$ for some $\varepsilon>0$, where $\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})-\widehat{\theta}^{0}_{n}(t,a)$ can be expressed as
$\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})-\widehat{\theta}^{0}_{n}(t,a)$ =$=\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})-\theta^{0}(t,a)\right\}-\left\{\widehat{\theta}^{0}_{n}(t,a)-\theta^{0}(t,a)\right\}$

### v2

@@ D.2.3 Lemmata for the source-site estimator
by Bayes’s rule. Furthermore, by Duhamel equation in Gill and Johansen [22] used in Lemma D.5 again, $\mathbb{P}[\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P}_{\infty})]-\theta^{0}(t,a)$ equals
$\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}S^{k}_{\infty}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})}{S^{k}_{\infty}(y\mid a,\mathbf{X})}$
×$\qquad\times\left\{\frac{\omega^{k,0}_{\infty}(\mathbf{X})G^{k}(y\mid a,\mathbf{X})\pi^{k}(a\mid\mathbf{X})}{\omega^{k,0}(\mathbf{X})G^{k}_{\infty}(y\mid a,\mathbf{X})\pi^{k}_{\infty}(a\mid\mathbf{X})}-1\right\}(\Lambda^{k}_{\infty}-\Lambda^{k})(dy\mid a,\mathbf{X})\bigg{]}.$ (8)
@@
By (D.2.3) with $\pi_{\infty}^{k}=\pi^{k}$, $\omega^{k,0}_{\infty}=\omega^{k,0}$, $G_{\infty}^{k}=G^{k}$, and $S_{\infty}^{k}=S^{k}$,
$\widehat{\theta}^{k,0}_{n}(t,a)-\theta^{0}(t,a)$ $=\mathbb{P}_{n}[\varphi^{*k,0}_{t,a}]+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}^{1/2}}{n}\mathbb{G}_{n}^{m}\left[\widehat{\varphi}^{k,0}_{n,m,t,a}-\varphi^{k,0}_{t,a}\right]$
$\qquad+\frac{1}{M}\sum_{m=1}^{M}\frac{Mn_{m}}{n}\mathbb{P}\left[\widehat{\varphi}^{k,0}_{t,a}-\theta^{0}(t,a)\right].$
By Conditions D.6 and D.7, the second summand on the right-hand-side is $o_{p}(n^{-1/2})$ by Lemma D.10. By Lemma D.11, $\mathbb{P}[\widehat{\varphi}^{k,0}_{t,a}]-\theta^{0}(t,a)$ equals
$\mathbb{E}\bigg{[}\frac{q^{0}(\mathbf{X})}{\mathbb{P}(R=0)}\widehat{S}^{k}_{m}(t\mid a,\mathbf{X})\int_{0}^{t}\frac{S^{k}(y-\mid a,\mathbf{X})}{\widehat{S}^{k}_{m}(y\mid a,\mathbf{X})}$
×$\qquad\times\left\{\frac{\widehat{\omega}^{k,0}_{m}(\mathbf{X})G^{k}(y\mid a,\mathbf{X})\pi^{k}(a\mid\mathbf{X})}{\omega^{k,0}(\mathbf{X})\widehat{G}^{k}_{m}(y\mid a,\mathbf{X})\widehat{\pi}^{k}_{m}(a\mid\mathbf{X})}-1\right\}(\widehat{\Lambda}^{k}_{m}-\Lambda^{k})(dy\mid a,\mathbf{X})\bigg{]}.$
By a similar decomposition in the proof of Theorem 2.8 in Section D.1.4 using Duhamel equation in Gill and Johansen [22], and by notation in Condition D.8, we find that the above bias term can be bounded by $\eta^{2}\{r_{n,t,a,1}^{k}+r_{n,t,a,2}^{k}+r_{n,t,a,3}^{k}\}$ over $m$. Since $M^{-1}\sum_{m=1}^{M}n^{-1}Mn_{m}\leq 2$, we have
@@
by Condition D.8. This established the pointwise RAL property: $\widehat{\theta}^{k,0}_{n}(t,a)=\theta^{0}(t,a)+\mathbb{P}_{n}(\varphi^{*k,0}_{t,a})+o_{p}(n^{-1/2})$. Since $\varphi^{*k,0}_{t,a}$ is uniformly bounded, $\mathbb{P}\{(\varphi^{*k,0}_{t,a})^{2}\}<\infty$ and since $\mathbb{P}\{\varphi^{*k,0}_{t,a}\}=0$, it follows that
@@
For the uniform RAL, by conditions D.6, D.7 and D.8 and Lemma D.10,
@@
which is $o_{p}(n^{-1/2})$ by Condition D.10. Therefore, $\sup_{u\in[0,t]}\left|\widehat{\theta}^{k}_{n}(u,a)-\theta^{0}(u,a)-\mathbb{P}_{n}(\varphi^{*k,0}_{u,a})\right|=o_{p}(n^{-1/2})$. Since $\{\varphi^{*k,0}_{u,a}:u\in[0,t]\}$ is a uniformly bounded $\mathbb{P}$-Donsker class by Lemma D.9, $\{n^{1/2}\mathbb{P}_{n}\{\varphi^{*k,0}_{u,a}\}:u\in[0,t]\}$ converges weakly to a tight mean-zero Gaussian process with covariance $(u,v)\mapsto\mathbb{P}(\varphi^{*k,0}_{u,a}\varphi^{*k,0}_{v,a})$. ∎
@@
If we only need the consistency of $\widehat{\theta}_{n}^{k}(t,a)$, then condition $\pi_{\infty}^{k}=\pi^{k}$, $\omega^{k,0}_{\infty}=\omega^{k,0}$, $G_{\infty}^{k}=G^{k}$, and $S_{\infty}^{k}=S^{k}$ can be replaced by the following statement: For $\mathbb{P}$-almost all $\mathbf{x}$, there exist measurable sets $\mathcal{S}^{k}_{x},\mathcal{G}^{k}_{x}\subseteq[0,t]$ such that $\mathcal{S}^{k}_{x}\cup\mathcal{G}^{k}_{x}=[0,t]$ and $\Lambda^{k}(u\mid a,\mathbf{x})=\Lambda_{\infty}^{k}(u\mid a,\mathbf{x})$ for all $u\in S^{k}_{x}$ and $G(u\mid a,\mathbf{x})=G^{k}_{\infty}(u\mid a,\mathbf{x})$ for all $u\in\mathcal{G}^{k}_{x}$. In addition, if $\mathcal{S}^{k}_{x}$ is a strict subset of $[0,t]$, then $\pi^{k}(a\mid\mathbf{x})=\pi^{k}_{\infty}(a\mid\mathbf{x})$ and $\omega^{k,0}(\mathbf{x})=\omega^{k,0}_{\infty}(\mathbf{x})$ as well. Then, $\widehat{\theta}_{n}^{k}(t,a)$ is consistent if Conditions D.6 and D.7 hold, and it is uniform consistent if Condition D.9 also holds. This statement could be interpreted as that at a given time $t$, if either (i) the conditional survival model $S^{k}$; or (ii) all other nuisance functions $G^{k}$, $\pi^{k}$ and $\omega^{k,0}$ are correctly specified (with other conditions above), $\widehat{\theta}_{n}^{k}(t,a)$ is consistent.
The proof sketch of Remark D.12 is similar to that under Remark D.6 via examining the bias term (D.2.3), which is omitted here.
@@ D.3 Theory for the federated estimator
Let the set of all source site indices be $\mathcal{S}={1,\dots,K-1}$. We then define the oracle selection space for $\boldsymbol{\eta}_{t,a}$, and the corresponding weight space as:
$\mathcal{S}^{*}_{t,a}=\{k\in\mathcal{S}:\theta^{k}(t,a)=\theta^{0}(t,a)\},~{}~{}\text{and}~{}~{}\mathbb{R}^{S^{*}_{t,a}}=\{\boldsymbol{\eta}_{t,a}\in\mathbb{R}^{K-1}:\eta^{j}_{t,a}=0,\forall j\not\in\mathcal{S}^{*}_{t,a}\},$
respectively.
The space $\mathcal{S}^{*}_{t,a}$ is both time- and treatment-varying, indicating that a source site may not consistently be useful or unhelpful across different time points or treatments. However, it offers the advantage of increased flexibility and adaptivity, allowing for more effective borrowing of information at different points along the survival functions. Based on the theory presented in Section D.2, for $k\in\mathcal{S}^{*}_{t,a}$, the site-specific estimator $\widehat{\theta}^{k,0}_{n}(t,a)$ is consistent for $\theta^{0}(t,a)$ for any given $t\in[0,\tau]$ and $a\in\{0,1\}$.
We begin by assuming fixed $\boldsymbol{\eta}_{t,a}=(\eta_{t,a}^{0},\eta_{t,a}^{1},\dots,\eta_{t,a}^{K-1})$. We invoke Lemmata 4 and 5 in Han et al. [25], which state that the proposed adaptive estimation for $\eta_{t,a}^{k}$ as shown in (2) allows for (i) the recovery of the optimal $\bar{\eta}^{k}_{t,a}$ by the estimator $\widehat{\eta}^{k}_{t,a}$, and (ii) the uncertainty induced by $\widehat{\eta}^{k}_{t,a}$ is negligible when estimating $\theta^{0}(t,a)$. We require regularity Conditions D.6, D.7 and D.8 for the pointwise convergence result in Theorem D.7 hold. Let us denote the federated estimator by plugging-in the fixed $\boldsymbol{\eta}_{t,a}$ as
@@
Recall that notation $\mathcal{H}_{t,a}$ defined in (1):
$\mathcal{H}_{t,a}(\mathcal{O};S,G)=\frac{\mathbb{I}(Y\leq t,\delta=1)}{S(Y\mid a,\mathbf{X})G(Y\mid a,\mathbf{X})}-\int_{0}^{t\wedge Y}\frac{\Lambda(du\mid a,\mathbf{X})}{S(u\mid a,\mathbf{X})G(u\mid a,\mathbf{X})}.$
Let us then write
$\xi^{0,(1)}(\mathcal{O})$ $=S^{0}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\pi^{0}(a\mid\mathbf{X})}\mathcal{H}_{t,a}(\mathcal{O};S^{0},\Lambda^{0},G^{0}),$
$\xi^{k,0,(1)}(\mathcal{O})$ $=\omega^{k,0}(\mathbf{X})S^{k}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}\mathcal{H}_{t,a}(\mathcal{O};S^{k},\Lambda^{k},G^{k}),$
@@
$\widehat{\theta}^{\text{fed}}_{n}(t,a;\boldsymbol{\eta}_{t,a})-\theta^{0}(t,a)$
$=\left(1-\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\right)\left\{\widehat{\theta}^{0}_{n}(t,a)-\theta^{0}(t,a)\right\}+\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\left\{\widehat{\theta}^{k,0}_{n}(t,a)-\theta^{0}(t,a)\right\}$
@@
The asymptotic variance of $\widehat{\theta}^{\text{fed}}_{n}(t,a;\boldsymbol{\eta}_{t,a})$ equals the variance of the influence function of (D.3). Let us denote it as $\mathcal{V}_{t,a}=\mathcal{V}_{t,a}(\boldsymbol{\eta}_{t,a})$. We highlight its dependence to the federated weights vector $\boldsymbol{\eta}_{t,a}$ here because in the below (11), we consider an optimization program for deriving the weights based on minimizing the (estimated) asymptotic variance.
Under the assumption of i.i.d. participants within each site, we have
$\mathcal{V}_{t,a}$ $=\left(1-\sum_{k\in\mathcal{S}}\eta^{k}_{t,a}\right)^{2}\frac{\mathbb{V}\{\xi^{0,(2)}(\mathcal{O}_{i})-\xi^{0,(1)}(\mathcal{O}_{i})\mid R_{i}=0\}}{\mathbb{P}(R_{i}=0)}$
@@
$\sqrt{n}\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a;\boldsymbol{\eta}_{t,a})-\theta^{0}(t,a)\right\}\to_{d}\mathcal{N}(0,\mathcal{V}_{t,a}).$
@@
$\bar{\boldsymbol{\eta}}_{t,a}=\underset{\eta^{k}_{t,a}=0,\forall k\not\in\mathcal{S}^{*}_{t,a}}{\text{arg min}}\mathcal{V}_{t,a}(\boldsymbol{\eta}_{t,a}).$ (11)
By leveraging Lemmata 4 and 5 from Han et al. [25], we can recover the optimal weights $\bar{\boldsymbol{\eta}}_{t,a}$ with negligible uncertainty for estimating $\theta^{0}(t,a)$ if we estimate ${\boldsymbol{\eta}}_{t,a}$ using (2), akin to adaptive Lasso [57, 18]. The consistency of $\widehat{\mathcal{V}}_{t,a}=\widehat{\mathcal{V}}_{t,a}(\widehat{\boldsymbol{\eta}}_{t,a})$ follows when we can effectively approximate ${\mathcal{V}}_{t,a}(\bar{\boldsymbol{\eta}}_{t,a})$ with $\widehat{\mathcal{V}}_{t,a}$. Thus,
$\sqrt{n/\widehat{\mathcal{V}}_{t,a}}\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a)-\theta^{0}(t,a)\right\}\to_{d}\mathcal{N}(0,1).$
@@
To establish that the asymptotic variance of $\widehat{\theta}^{\text{fed}}_{n}(t,a;\widehat{\boldsymbol{\eta}}_{t,a})$ is strictly smaller than that of the estimator based solely on the target data $\widehat{\theta}^{0}_{n}(t,a)$, we adopt Proposition 1 in Han et al. [25] with a modified informative source condition (modified Assumption 3(b) in Han et al. [25]).
Specifically, for each source site $s\in\mathcal{S}^{*}_{t,a}$, we define $\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})$ a federated estimator where $\eta^{s}_{t,a}$ is the optimal ensemble weight of site $s$ if we only consider target site and this source site $s$ for the federation. Then, the modified informative source condition is given as
$\left|\text{Cov}\left[\sqrt{n}\widehat{\theta}^{0}_{n}(t,a),\sqrt{n}\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})-\widehat{\theta}^{0}_{n}(t,a)\right\}\right]\right|\geq\varepsilon,$
for some $\varepsilon>0$, where $\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})-\widehat{\theta}^{0}_{n}(t,a)$ can be expressed as
$\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})-\widehat{\theta}^{0}_{n}(t,a)$
=$=\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a;\eta^{s}_{t,a})-\theta^{0}(t,a)\right\}-\left\{\widehat{\theta}^{0}_{n}(t,a)-\theta^{0}(t,a)\right\}$